A body is projected with a velocity 60m/s at 30 to horizontal. Its velocity after 3 sec?

Grade:11

1 Answers

Aman gohel
6 years ago
X component,Ux=ucos30=60cos30. =30√3m/s. Y component,Uy=usin30=60*0.5. =30m/sVelocity after 3 s, Vy=Uy-gt =>Vy=30-10*3 =>Vy=0Resultant velocity,R=√((30√3)^2+0^2)[According to triangle`s law of vector addition]=>R=52m/s(Ans)